OpenStudy (anonymous):

25 = 100(0.95)^(n-1)

OpenStudy (anonymous):

\[25 = 100(0.95)^{n-1}\]

OpenStudy (anonymous):

\[.25=.95^{n-1}\] \[n-1=\frac{\ln(.25)}{\ln(.95)}\] then a calculator

OpenStudy (anonymous):

I don't understand... What is \[(0.95)^{n-1}\]

OpenStudy (anonymous):

it is what you wrote

OpenStudy (anonymous):

It's not 0.95n -0.95?

OpenStudy (anonymous):

so for example if \(n=4\) you would have \((.95)^{4-1}=.95^3=0.857375\)

OpenStudy (anonymous):

It's not 25 = 100(0.95n -0.95)

OpenStudy (anonymous):

no not if the \(n-1\) is in the exponent

OpenStudy (anonymous):

I am going to remake this question satelite.

OpenStudy (anonymous):

ok